Solve Problems Using Linear Systems
14. The perimeter of a pool table is about 7.8 m. Four times the length equals nine times the width. What are the dimensions of the table, in metres?
Can you please help me to solve this problem? thank you very much!!!!

Answer by checkley77(12844)

4L=9W OR L=9W/4
2L+2W=7.8
2(9W/4)+2W=7.8
18W/4+2W=7.8
(18W+4*2W)/4=7.8
(18W+8W)/4=7.8
26W=4*7.8
26W=31.2
W=31.2/26
W=1.2 M. IS THE WIDTH OF THE POOL TABLE.
L=9*1.2/4
L=10.8/4
L=2.7M. IS THE LENGTH OF THE POOL TABLE.
PROOF:
2*2.7+2*1.2=7.8
5.4+2.4=7.8
7.8=7.8